Movie Twitter Activity Receipts The Devil Inside 219509 14763 The Dictator 6405 5796 Paranormal Activity 3...

Movie

Twitter Activity

Receipts

The Devil Inside

219509

14763

The Dictator

6405

5796

Paranormal Activity 3

165128

15829

The Hunger Games

579288

36871

Bridesmaids

6564

8995

Red Tails

11104

7477

Act of Valor

9152

8054

1) What is the R-squared value and interpret it’s meaning?

0.97597921

2) What is the correlation coefficient and interpret it’s meaning?

0.9879166

3) Predict the receipts for a movie that has a Twitter activity of 100,000 using the linear regression formula generated after you run the regression analysis in Excel (just like we did in the lecture notes):

Receipts = 7,310.82

4) At the 95% significance level, is there significant evidence of a relationship between Twitter activity and receipts?

5) Do you accept or reject the null hypothesis?

6) Based upon your answers above, provide us with a paragraph (3-4 sentence) conclusion that would be appropriate to explain the results to a senior executive that isn’t as familiar with statistics as you are!
Answer #1
SUMMARY OUTPUT
Regression Statistics
Multiple R 0.987917
R Square 0.975979
Adjusted R Square 0.971175
Standard Error 1830.236
Observations 7
ANOVA
df SS MS F Significance F
Regression 1 6.81E+08 6.81E+08 203.153 3.06E-05
Residual 5 16748821 3349764
Total 6 6.97E+08
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 6808.105 854.9682 7.962991 0.000504 4610.339 9005.87 4610.339 9005.87
x 0.050272 0.003527 14.25318 3.06E-05 0.041205 0.059338 0.041205 0.059338

so, regression line is   Ŷ =   6808.105 +   0.050272 *x

3)

Predicted Y at X=   100000   is                  
Ŷ =   6808.10470   +   0.050272   *   100000   =   11835.259

4)

Ho:   ß1=   0          
H1:   ß1╪   0          
n=   7              
alpha =   0.05              
estimated std error of slope =Se(ß1) = Se/√Sxx =    1830.236   /√   269273194950.00   =   0.0035
                  
t stat = estimated slope/std error =ß1 /Se(ß1) =    0.0503   /   0.0035   =   14.2532
                  
Degree of freedom ,df = n-2=   5              
p-value =    0.0000              
decison :    p-value<α , reject Ho  

SO there IS significant evidence of a relationship between Twitter activity and receipts

           
5)

WE will reject the null hypothesis concluding that there is linear relationship

6)


from above ,

we can say that for a movie, the twitter activity and receipts are linearly related. As the twitter activity increases the receipt will also increase and vice versa

the twitter activity explain the response of receipts.